OpenStudy (anonymous):

find the exact value of sin^-1(-sqrt3/2) without using a calculator. answer in radians

OpenStudy (anonymous):

\[\sin^{-1} (-\sqrt{3}/2)\] \[-\sin^{-1} (\sqrt{3}/2)\] \[-\pi/3\]

OpenStudy (anonymous):

becuase \[\sin (\pi/3)=\sqrt{3}/2\]

OpenStudy (anonymous):

and \[\sin^{-1} (-x)=-\sin^{-1} (x)\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

u are welcome...

OpenStudy (anonymous):

any other doubts...

OpenStudy (anonymous):

yeh one more can you show tan^-1(11/7)-tan^-1(2/9)=pi/4 without using a calculator?

OpenStudy (anonymous):

ya...

OpenStudy (anonymous):

apply the formula... \[\tan^{-1} (x)-\tan^{-1} (y)=\tan^{-1} ((x-y)/(1+xy))\]

OpenStudy (anonymous):

x=11/7 and y=2/9

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

its ok